please help me solve this equation: x/x+1-2=3/x-3
Answer by persian52(161)

(x)/(x)+1-2=(3)/(x)-3
Find the LCD (least common denominator) of (x)/(x)+1-2+(3)/(x)-3.
Least common denominator: x
Multiply each term in the equation by x in order to remove all the denominators from the equation.
(x)/(x)*x+1*x-2*x=(3)/(x)*x-3*x
Simplify the left-hand side of the equation by canceling the common terms.
0=(3)/(x)*x-3*x
Simplify the right-hand side of the equation by simplifying each term.
0=-3x+3
Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
-3x+3=0
Since 3 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3 from both sides.
-3x=-3
Divide each term in the equation by -3.
-(3x)/(-3)=-(3)/(-3)
Simplify the left-hand side of the equation by canceling the common terms.
x=-(3)/(-3)
Simplify the right-hand side of the equation by simplifying each term.
Answer: x=1